OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (-3, 2) and (2, 1). y + 3 = -5(x - 2) y - 2 = -5(x + 3) y + 3 = -one fifth(x - 2) y - 2 = -one fifth(x + 3)

OpenStudy (anonymous):

@jagr2713

jagr2713 (jagr2713):

y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −(x − 2) y − 2 = −(x + 3) is this it?

OpenStudy (anonymous):

yup

jagr2713 (jagr2713):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1)? The equation is of the form y - y1 = m(x - x1) (x1, y1) is either of the given points; "m" is the slope. Find the slope. m = (y2-y1)/(x2-x1) ....= (1-2)/(2+3) = -1/5 However, this slope is not in any of the options, and I checked my work, so either you mistyped something, maybe one of the points, or the people making the book made a mistake. Plug in either point for y1 and x1 in the equation. Apparently they used (-3, 2). y - 2 = -(1/5)(x - (-3)) y - 2 = -(1/5)(x + 3) so its D